# Transition temperature of homogeneous weakly interacting Bose gas in self-consistent Popov approximation

### v1

@@ Abstract
Using the Cornwall–Jackiw–Tomboulis effective action framework in conjunction with variational perturbation theory, this study investigates the relative shift in the transition temperature of a homogeneous repulsive weakly interacting Bose gas compared to that of an ideal Bose gas. Employing both one-loop and self-consistent Popov approximations, we derive the universal form of the relative shift in the transition temperature, which is proportional to the $s$-wave scattering length. The results demonstrate excellent agreement with those obtained through precise Monte Carlo simulations. Furthermore, the zero-point energy and various thermodynamic quantities are also analyzed in the condensed phase.
@@ I Introduction
where the gas parameter $\alpha_{s}=\rho a_{s}^{3}$ is defined in terms of $\rho$, the particle number density. The determination of the constants $a$ and $c$ has been the subject of prolonged debate and controversy. Most studies suggest $a=1/3$ [9, 10, 11, 12], though some results indicate $a=1/6$ [13, 14]. Similarly, the value of the constant $c$ has been contentious: while some calculations yield negative values [9, 15], the majority report positive values over a broad range [10, 11, 16].
@@
To study the dilute Bose gas at finite temperatures, most authors have utilized the Hartree-Fock-Bogoliubov (HFB) theory, as documented in the literature [5, 17, 18]. However, the HFB theory presents certain disadvantages, including issues related to conservation laws and the presence of a gap in the energy spectrum of excitations, which arise due to the anomalous density, as discussed in Ref. [19]. To address these challenges, several approximations have been proposed, such as the Popov, Beliaev-Popov, and $T$-matrix approximations [20, 21], with the Popov approximation being particularly prevalent. Nevertheless, the Popov approximation results in a first-order phase transition in the condensation phase transition, which is an unphysical outcome. To overcome this inconsistency, Yakulov and collaborators [22] introduced a framework involving Lagrange multipliers with two chemical potentials. In their approach, one chemical potential primarily resolves the conservation problem, while the other ensures the energy spectrum remains gapless. Despite this refinement, their results indicated that the transition temperature for a weakly interacting Bose gas coincides with that of an ideal Bose gas. A similar issue was encountered when Haugset et. al. [23] applied the Cornwall–Jackiw–Tomboulis (CJT) effective action approach in the one-loop approximation to investigate the dilute Bose gas. Fortunately, this conclusion was subsequently disproven and addressed by Kleinert et. al. [13] through the application of variational perturbation theory. While the relative shift in the transition temperature was found to deviate from zero, the authors clarified that their result, characterized by a square-root behavior, was incorrect. In the present study, we integrate the CJT effective action approach within the one-loop approximation with variational perturbation theory in the Popov approximation. This combination aims to refine and extend the findings in Refs. [13, 23], particularly concerning the relative shift in transition temperature and the thermodynamic quantities of a homogeneous repulsive weakly interacting Bose gas.
@@
This paper is organised as follows. In Section II focus on finding the relative shift of transition temperature and thermodynamic quantities of a homogeneous dilute weakly interacting Bose gas by means of the CJT effective action approach at finite temperature. Finally, we present the conclusion and future outlook in Section III.
II Transition temperature and thermodynamic quantities of a homogeneous dilute Bose gas
@@ II.1 Self-consistent Popov approximation
wherein $\mu$ is the chemical potential. The field operator $\psi(\vec{r},t)$ depends on both the coordinate $\vec{r}$ and time $t$. The interatomic interaction potential between the atoms can be chosen as the hard-sphere model. In the Born approximation, the strength of the interaction between pairwise atoms is determined via the $s$-wave scattering length $a_{s}$ as $g=4\pi\hbar^{2}a_{s}/m$. Now, thermodynamic stability requires that $g>0$, i.e., the boson interactions are repulsive [5, 17]. In the tree-approximation, the expectation value $\psi_{0}$ of the field operator is independent of both coordinate and time. Therefore, the GP potential can be read off from the Lagrangian density (3),
@@
$V_{\beta}=-\mu\left|\psi_{0}\right|^{2}+\frac{g}{2}\left|\psi_{0}\right|^{4}+\frac{1}{2}\int_{\beta}\frac{d^{3}\vec{k}}{(2\pi)^{3}}E(k)+\frac{1}{\beta}\int\frac{d^{3}\vec{k}}{(2\pi)^{3}}\ln\left[1-e^{-\beta E(k)}\right].$ (15)
By extremizing the CJT effective potential (15) with respect to $\psi_{0}$ one arrives at the gap equation in the one-loop approximation
@@
$+\int\frac{d^{3}\vec{k}}{(2\pi)^{3}}\frac{2g\left[\varepsilon_{k}-\mu+2g\left|\psi_{0}\right|^{2}\right]-g^{2}\left|\psi_{0}\right|^{2}}{E(k)\left[e^{\beta E(k)}-1\right]}.$ (16)
@@
$\rho=\frac{\partial P}{\partial\mu}.$ (19)
@@
$\rho-\rho_{0}=\int\frac{d^{3}\vec{k}}{(2\pi)^{3}}\frac{\varepsilon_{k}+g\rho_{0}}{\sqrt{\varepsilon_{k}^{2}+2g\rho_{0}\varepsilon_{k}}}\left[\frac{1}{2}+\frac{1}{e^{\beta\sqrt{\varepsilon_{k}^{2}+2g\rho_{0}\varepsilon_{k}}}-1}\right].$ (21)
@@
To ensure self-consistency, the variational perturbative method introduced by Kleinert et.al. [13] is employed. This method involves introducing an expansion parameter $\eta$ and a variational parameter $M$ defined by the relation $\mu=M+r\eta$ where $r=(\mu-M)/\eta$. These parameters result in a trial CJT effective potential $V_{\beta}^{\rm(trial)}$, replacing $V_{\beta}$ in Eq. (15). The optimization of $V_{\beta}^{\rm(trial)}$ is performed by requiring that $\partial V_{\beta}^{\rm(trial)}/\partial M=0$ at $M=M^{\rm(opt)}$ while maintaining a fixed temperature. This yields
@@
This condition determines $M=M^{\rm(opt)}=gn$. As a result, the self-consistent pressure is obtained, maintaining the same form as Eq. (18), with the condensed density $\rho_{0}$ substituted by the total particle density $\rho$. For simplicity the self-consistent pressure is still denoted by ${\cal P}$
${\cal P}$ $=$ $\frac{\left[M^{\rm(opt)}\right]^{2}}{2g}-\frac{1}{2}\int\frac{d^{3}\vec{k}}{(2\pi)^{3}}\sqrt{\varepsilon_{k}^{2}+2gn\varepsilon_{k}}-\frac{1}{\beta}\int\frac{d^{3}\vec{k}}{(2\pi)^{3}}\ln\left[1-e^{-\beta\sqrt{\varepsilon_{k}^{2}+2gn\varepsilon_{k}}}\right]$ (23)
@@
A comparison between Eqs. (21) and (24) reaffirms that the results derived under the self-consistent Popov approximation exhibit the same forms as those within the conventional Popov approximation, with the total particle density $\rho$ replacing the condensed density $\rho_{0}$.
@@ II.2 Transition temperature
where $\Gamma(x)$ is the gamma function, $\Omega_{d}=2\pi^{d/2}/\Gamma(d/2)$ is the surface area of a $d-$dimensional sphere. Employing (26) the integral (25) is calculated
@@
$I=$ $\int\frac{d^{3}\vec{k}}{(2\pi)^{3}}\frac{\varepsilon_{k}+g\rho}{\sqrt{\varepsilon_{k}^{2}+2g\rho\varepsilon_{k}}}\frac{1}{e^{\beta\sqrt{\varepsilon_{k}^{2}+2g\rho\varepsilon_{k}}}-1}=\frac{(2m)^{3/2}}{2\pi^{2}\hbar^{3}}$ (29)
×$\times\int_{0}^{\infty}dq\frac{q^{2}(q^{2}+\beta g\rho)}{\sqrt{q^{2}(q^{2}+2\beta g\rho)}}\frac{1}{e^{\sqrt{q^{2}(q^{2}+2\beta g\rho)}}-1}.$
@@
where $\rho_{\rm ex}$ is defined as the non-condensed density. It is notable that the first term on the right-hand side of Eq. (31) corresponds to quantum fluctuations, which reduce the condensed density due to interaction-induced quantum effects. This contribution was initially identified by Bogoliubov [31] and subsequently corroborated by other studies, such as Refs. [6, 30].
At the transition temperature, the condensed density vanishes ($\rho_{0}=0$). From Eq. (31), the relative shift in the transition temperature, to the lowest order of the $s$-wave scattering length, is given by
@@
$c=-\frac{4\zeta(1/2)}{3\zeta(3/2)^{1/3}}\approx 1.413.$ (33)
This value exhibits a high degree of agreement with our result when compared to those obtained from precise Monte Carlo simulations [34, 35] $c=1.29\pm 0.05$. Furthermore, it is in excellent concordance with the theoretical result derived using the nonperturbative linear $\delta$-expansion method [36] $c=1.32\pm 0.02$ and aligns closely with other finding in mathematical physics [12] $c=1.49$.
II.3 Zero-point energy and thermodynamic quantities
In this section, we examine the temperature dependence of the zero-point energy and thermodynamic quantities in the condensed phase of the gas. The initial focus is placed on the chemical potential. Inversion of Eq. (20) one has
@@ II.3 Zero-point energy and thermodynamic quantities
By employing Eqs. (26) and following a procedure analogous to that used for Eq. (29), the chemical potential can be determined and expressed as a function of temperature
$\mu=g\rho\left(1+\frac{32}{3\sqrt{\pi}}\alpha_{s}^{1/2}\right)-\frac{24\pi\hbar^{2}\zeta(1/2)\rho a_{s}^{2}}{m\lambda_{B}}+\frac{4\pi\hbar^{2}\zeta(3/2)a_{s}}{m\lambda_{B}^{3}}.$ (35)
@@
${\cal P}=\frac{1}{2}g\rho^{2}\left(1+\frac{128}{15\sqrt{\pi}}\alpha_{s}^{1/2}\right)+\frac{4\pi\hbar^{2}\zeta(3/2)}{m\lambda_{B}^{3}}\rho a_{s}-\frac{6\pi\hbar^{2}\zeta(3/2)^{2}a_{s}}{m\lambda_{B}^{6}}+\frac{2\pi\hbar^{2}\zeta(5/2)}{m\lambda_{B}^{5}}.$ (41)
@@
${\cal E}=\frac{1}{2}g\rho^{2}\left(1+\frac{128}{15\sqrt{\pi}}\alpha_{s}^{1/2}\right)-\frac{4\pi\hbar^{2}\zeta(3/2)\rho a_{s}}{m\lambda_{B}^{3}}+\frac{2\pi\hbar^{2}\zeta(3/2)^{2}a_{s}}{m\lambda_{B}^{6}}-\frac{2\pi\hbar^{2}\zeta(5/2)}{m\lambda_{B}^{5}}.$ (43)
@@
III Conclusion and outlook

### v2

# Transition temperature and thermodynamic properties of homogeneous weakly interacting Bose gas in self-consistent Popov approximation

@@ Abstract
This study utilizes the Cornwall-Jackiw-Tomboulis effective action approach combined with variational perturbation theory to investigate the relative shift in the transition temperature of a homogeneous, repulsive, weakly interacting Bose gas compared to that of an ideal Bose gas. By applying both the one-loop and self-consistent Popov approximations, the universal form of the relative shift in the transition temperature is derived, demonstrating its proportionality to the $s$-wave scattering length. The results exhibit excellent agreement with those obtained from precise Monte Carlo simulations. Furthermore, the zero-point energy and various thermodynamic properties are examined in both the condensed and normal phases. A comparison with experimental data reveals an excellent agreement, further validating the findings.
@@ I Introduction
where the gas parameter $\alpha_{s}=\rho a_{s}^{3}$ is defined in terms of $\rho$, the particle number density and the scattering length $a_{s}$. The determination of the constants $a$ and $c$ has been the subject of prolonged debate and controversy. Most studies suggest $a=1/3$ [9, 10, 11, 12] (i.e. the shift is proportional to $a_{s}$ and $\rho^{1/3}$), though some results indicate $a=1/6$ [13, 14, 15] (i.e. the shift is proportional to $a_{s}^{1/2}$ and $\rho^{1/6}$). Similarly, the value of the constant $c$ has been contentious: while some calculations yield negative values [9, 16], the majority report positive values over a broad range [10, 11, 17]. A remarkably high value of $c=3.059$ was reported in Ref. [18], while an even greater value of $c=4.66$ was documented in Ref. [19]. Furthermore, experimental data have indicated a significant value of $c=5.1\pm 0.9$ in Ref. [20].
Most recent investigations [21] have suggested that, at low densities, the relative shift exhibits a dependence on $\rho^{2/3}$, whereas alternative analyses [22] propose a scaling with $\rho^{1/6}$ and $a_{s}^{1/2}$. The numerical computation results are not consistent. Specifically, Grüter et.al. obtained $c=0.34\pm 0.06$ in Ref. [10], whereas the authors of Refs. [23, 24] reported $c=1.29\pm 0.05$.
@@
To study the dilute Bose gas at finite temperatures, most authors have utilized the Hartree-Fock-Bogoliubov (HFB) theory, as documented in the literature [5, 25]. However, the HFB theory presents certain disadvantages, including issues related to conservation laws and the presence of a gap in the energy spectrum of excitations, which arise due to the anomalous density, as discussed in Ref. [26]. To address these challenges, several approximations have been proposed, such as the Popov, Beliaev-Popov, and $T$-matrix approximations [27, 28], with the Popov approximation being particularly prevalent. Nevertheless, the Popov approximation results in a first-order phase transition in the condensation phase transition, which is an unphysical outcome. To overcome this inconsistency, Yakulov and collaborators [29] introduced a framework involving Lagrange multipliers with two chemical potentials. In their approach, one chemical potential primarily resolves the conservation problem, while the other ensures the energy spectrum remains gapless. Despite this refinement, their results indicated that the transition temperature for a weakly interacting Bose gas coincides with that of an ideal Bose gas. A similar issue was encountered when Haugset et. al. [30] applied the Cornwall–Jackiw–Tomboulis (CJT) effective action approach in the one-loop approximation to investigate the dilute Bose gas. Fortunately, this conclusion was subsequently disproven and addressed by Kleinert et. al. [13] through the application of variational perturbation theory. While the relative shift in the transition temperature was found to deviate from zero, the authors clarified that their result, characterized by a square-root behavior, was incorrect. In the present study, we integrate the CJT effective action approach within the one-loop approximation with variational perturbation theory in the Popov approximation. This combination aims to refine and extend the findings in Refs. [13, 30], particularly concerning the relative shift in transition temperature and the thermodynamic quantities of a homogeneous repulsive weakly interacting Bose gas. Indeed, thermodynamic properties of the interacting Bose gas are investigated in both condensed and normal phases.
@@
This paper is organised as follows. In Section II focus on finding the relative shift of transition temperature with respect to the ideal Bose gas by means of the CJT effective action approach at finite temperature. Section III devotes for thermodynamic quantities of a homogeneous dilute weakly interacting Bose gas . Finally, we present the conclusion and future outlook in Section IV.
II Transition temperature in the self-consistent Popov approximation
@@ II.1 Self-consistent Popov approximation
wherein $\mu$ is the chemical potential. The field operator $\psi(\vec{r},t)$ depends on both the coordinate $\vec{r}$ and time $t$. The interatomic interaction potential between the atoms can be chosen as the hard-sphere model. In the Born approximation, the strength of the interaction between pairwise atoms is determined via the $s$-wave scattering length $a_{s}$ as $g=4\pi\hbar^{2}a_{s}/m$. Now, thermodynamic stability requires that $g>0$, i.e., the boson interactions are repulsive [5]. In the tree-approximation, the expectation value $\psi_{0}$ of the field operator is independent of both coordinate and time. Therefore, the GP potential can be read off from the Lagrangian density (3),
@@
$V_{\beta}=$ $-\mu\left|\psi_{0}\right|^{2}+\frac{g}{2}\left|\psi_{0}\right|^{4}+\frac{1}{2}\int_{\beta}\frac{d^{3}\vec{k}}{(2\pi)^{3}}E(k)$ (15)
$+\frac{1}{\beta}\int\frac{d^{3}\vec{k}}{(2\pi)^{3}}\ln\left[1-e^{-\beta E(k)}\right],$
By extremizing the CJT effective potential (15) with respect to $\psi_{0}$ one arrives at the gap equation
@@
$+\int\frac{d^{3}\vec{k}}{(2\pi)^{3}}\frac{2g\left[\varepsilon_{k}-\mu+2g\left|\psi_{0}\right|^{2}\right]-g^{2}\left|\psi_{0}\right|^{2}}{E(k)\left[e^{\beta E(k)}-1\right]}=0.$ (16)
@@
$\rho=\frac{\partial{\cal P}}{\partial\mu}.$ (19)
@@
$\rho-\rho_{0}=\int\frac{d^{3}\vec{k}}{(2\pi)^{3}}\frac{\varepsilon_{k}+\mu}{\sqrt{\varepsilon_{k}^{2}+2\mu\varepsilon_{k}}}\left[\frac{1}{2}+\frac{1}{e^{\beta\sqrt{\varepsilon_{k}^{2}+2\mu\varepsilon_{k}}}-1}\right].$ (21)
@@
To ensure self-consistency, the variational perturbation method introduced by Kleinert et.al. [13] is employed. This method involves introducing a variational parameter $M$ defined by the relation $\mu=M+r\eta$ where $\eta$ is a counting loop expansion parameter and $r=(\mu-M)/\eta$. This parameter results in a trial CJT effective potential $V_{\beta}^{\rm(trial)}$, replacing $V_{\beta}$ in Eq. (15). The optimization of $V_{\beta}^{\rm(trial)}$ is performed by requiring that $\partial V_{\beta}^{\rm(trial)}/\partial M=0$ at $M=M^{\rm(opt)}$ while maintaining a fixed temperature. Reinserting $r$ into the obtained result and setting $\eta=1$ yields
@@
This condition determines $M=M^{\rm(opt)}=g\rho$. As a result, the self-consistent pressure is obtained, maintaining the same form as Eq. (18), with the condensed density $\rho_{0}$ substituted by the total particle density $\rho$. For simplicity the self-consistent pressure is still denoted by ${\cal P}$
${\cal P}$ $=$ $\frac{\left[M^{\rm(opt)}\right]^{2}}{2g}-\frac{1}{2}\int\frac{d^{3}\vec{k}}{(2\pi)^{3}}\sqrt{\varepsilon_{k}^{2}+2g\rho\varepsilon_{k}}-\frac{1}{\beta}\int\frac{d^{3}\vec{k}}{(2\pi)^{3}}\ln\left[1-e^{-\beta\sqrt{\varepsilon_{k}^{2}+2g\rho\varepsilon_{k}}}\right]$ (23)
@@
Equation (24) serves as a gap equation, governing the evolution of the system under varying parameters.
@@ II.2 Transition temperature
where $\Gamma(x)$ is the gamma function, $\Omega_{d}=2\pi^{d/2}/\Gamma(d/2)$ is the surface area of a $d-$dimensional sphere and $\Lambda$ is a renormalization scale that ensures the integral has the correct canonical dimension. Employing (26) the integral (25) is calculated
@@
$I$ $=$ $\int\frac{d^{3}\vec{k}}{(2\pi)^{3}}\frac{\varepsilon_{k}+g\rho}{\sqrt{\varepsilon_{k}^{2}+2g\rho\varepsilon_{k}}}\frac{1}{e^{\beta\sqrt{\varepsilon_{k}^{2}+2g\rho\varepsilon_{k}}}-1}$ (29)
[…]
@@
where $\rho_{\rm ex}$ is defined as the non-condensed density. It is notable that the first term on the right-hand side of Eq. (31) corresponds to quantum fluctuations, which reduce the condensed density due to interaction-induced quantum effects.
We now calculate the transition temperature. At the transition point, the condensed density vanishes ($\rho_{0}=0$). From Eq. (31), the relative shift in the transition temperature, to the lowest order of the $s$-wave scattering length, is given by
@@
$c=-\frac{4\zeta(1/2)}{3\zeta(3/2)^{1/3}}\approx 1.413\approx\sqrt{2},$ (33)
and $a=1/3$. These values exhibit a high degree of agreement with our result when compared to those obtained from precise Monte Carlo simulations [23, 24] $c=1.29\pm 0.05$. Furthermore, it is in excellent concordance with the theoretical result derived using the nonperturbative linear $\delta$-expansion method [41] $c=1.32\pm 0.02$. It also aligns closely with other finding in mathematical physics [12] $c=1.49$. Unfortunately, this result indicates a first-order phase transition.
III Zero-point energy and thermodynamic properties of the dilute Bose gas
III.1 Condensed phase
We now examine the temperature dependence of the zero-point energy and thermodynamic quantities in the condensed phase of the gas. The initial focus is placed on the chemical potential. Inversion of Eq. (20) one has
@@ III.1 Condensed phase
By employing Eqs. (26) and following a procedure analogous to that used for Eq. (29), the chemical potential can be determined and expressed as a function of temperature to the first order of the scattering length
$\mu=g\rho\left(1+\frac{32}{3\sqrt{\pi}}\alpha_{s}^{1/2}\right)+\frac{4\pi\hbar^{2}\zeta(3/2)a_{s}}{m\lambda_{B}^{3}}.$ (35)
@@
${\cal P}=$ $\frac{1}{2}g\rho^{2}\left(1+\frac{128}{15\sqrt{\pi}}\alpha_{s}^{1/2}\right)+\frac{4\pi\hbar^{2}\zeta(3/2)}{m\lambda_{B}^{3}}\rho a_{s}-\frac{6\pi\hbar^{2}\zeta(3/2)^{2}a_{s}}{m\lambda_{B}^{6}}$ (41)
$+\frac{2\pi\hbar^{2}\zeta(5/2)}{m\lambda_{B}^{5}}.$
@@
${\cal E}=$ $\frac{1}{2}g\rho^{2}\left(1+\frac{128}{15\sqrt{\pi}}\alpha_{s}^{1/2}\right)-\frac{4\pi\hbar^{2}\zeta(3/2)\rho a_{s}}{m\lambda_{B}^{3}}+\frac{2\pi\hbar^{2}\zeta(3/2)^{2}a_{s}}{m\lambda_{B}^{6}}$ (43)
$-\frac{2\pi\hbar^{2}\zeta(5/2)}{m\lambda_{B}^{5}}.$
@@
III.2 Normal phase
In the normal phase, that is above the transition temperature, the condensed fraction is zero, and consequently, all atoms remain in excited states. The dispersion relation, as given by Eq. (13), simplifies to
$E(k)=\varepsilon_{k}-\bar{\mu},$ (44)
and the propagator, represented by Eq. (12), reduces to
$D(k)=\frac{1}{\omega_{n}^{2}+E^{2}(k)}\left(\begin{array}[]{cc}\varepsilon_{k}-\bar{\mu}&\omega_{n}\\ -\omega_{n}&\varepsilon_{k}-\bar{\mu}\\ \end{array}\right).$ (45)
Herein we denote the chemical potential in the normal phase as $\bar{\mu}$, which is shifted $2g\rho$ relative to that of the ideal Bose gas. The ultraviolet divergence in the zero-point energy can be eliminated through the introduction of a counterterm, ensuring that it does not contribute to the pressure (as discussed in [30, 45]). The last term on the right-hand side of Equation (23) can be expressed using the polylogarithmic function, defined as ${\rm Li}_{n}(x)=\sum_{k=1}^{\infty}\frac{x^{k}}{n^{n}}$. The self-consistent pressure is then given by [51]
${\cal P}=g\rho^{2}+\frac{2\pi\hbar^{2}}{m\lambda_{B}^{5}}{\rm Li}_{5/2}\left[e^{\beta(\mu-2g\rho)}\right].$ (46)
Using (46) and (19) one can find the particle density
$\rho=\frac{{\rm Li}_{5/2}\left[e^{\beta(\mu-2g\rho)}\right]}{\lambda_{B}^{3}}.$ (47)
Inverting Eq. (47) yields the chemical potential
$\mu=2g\rho+k_{B}T\ln{\rm Li}_{3/2}^{-1}(\rho\lambda_{B}^{3}).$ (48)
Combining Eqs. (35) and (48) one has
$\mu=\left\{\begin{array}[]{ll}g\rho\left(1+\frac{32}{3\sqrt{\pi}}\alpha_{s}^{1/2}\right)+\frac{4\pi\hbar^{2}\zeta(3/2)a_{s}}{m\lambda_{B}^{3}},&\hbox{$T\leq T_{C}$;}\\ 2g\rho+k_{B}T\ln{\rm Li}_{3/2}^{-1}(\rho\lambda_{B}^{3}),&\hbox{$T\geq T_{C}$.}\end{array}\right.$ (51)
Figure 1: (Color online) The evolution of the chemical potential of the sodium 23 gas versus the reduced temperature. The black and red solid curves correspond to chemical potential of the interacting and ideal Bose gas. The blue open dots and red open triangles are the experimental data extracted from Ref. [51] for the fitting the density to the profile calculated at $T=280$ nK and 150 nK, respectively.
The evolution of the chemical potential with respect to the reduced temperature $T/T_{C}^{(0)}$, as defined in Eq. (51), is depicted by the black solid line in Fig. 1. The curve exhibits a continuous variation of the chemical potential across the transition point, further highlighting its non-monotonic behavior. Initially, in the condensed phase, the chemical potential increases monotonically with rising reduced temperature, starting from $g\rho$. It reaches a maximum value of $2g\rho$ (in fact the contribution from the quantum fluctuations is very small) at the critical point, after which it decreases as the temperature continues to increase. The red solid line, for reference, corresponds to the chemical potential of the ideal Bose gas.
We now compare our theoretical results with experimental data. Here we reference the experimental work of Mordini et al. [51], which investigated a BEC of sodium-23 isotopes confined in an Ioffe-Pritchard trap. In this experiment, parameters were chosen as $m=22.99u,\rho a_{s}^{3}=2\times 10^{-6}$. Using the local density approximation, the authors derived certain thermodynamic quantities for the corresponding homogeneous BEC. Experimental data for the chemical potential as a function of reduced temperature is plotted in Fig. 1. The blue open dots and red open triangles are the experimental data extracted from Ref. [51] for the fitting the density to the profile calculated at $T=280$ nK and 150 nK, respectively. These results reveal a strong agreement with our theoretical predictions. An important observation from Fig. 1 is that the experimental data confirm the non-monotonic behavior of the chemical potential in the vicinity of the transition point.
Similarly, the energy density above the transition temperature can be calculated. Utilizing Eq. (42), the expression is given by
${\cal E}=g\rho^{2}-\frac{2\pi\hbar^{2}}{m\lambda_{B}^{5}}{\rm Li}_{5/2}\left[e^{\beta(\mu-2g\rho)}\right].$ (52)
This result is consistent with the findings reported in Ref. [52].
To end this subsection, we emphasize that the contributions from quantum fluctuations have been ignored in the Eqs. (46), (48), (52).
IV Conclusion and outlook